Vector A⃗ has components Ax= 1.80 cm , Ay= 2.25 cm ; vector B⃗ has components Bx= 4.10 cm , By= -3.50 cm .

Find the x-component of the vector sum A⃗ +B⃗ .

Find the y-component of the vector sum A⃗ +B⃗ .

Find the magnitude of A⃗ +B⃗ .
Answer #1

Components of Vector A are:

Ax = 1.80 cm & Ay = 2.25 cm

Components of Vector B are:

Bx = 4.10 cm & By = -3.50 cm

Now Sum of x-component of vectors will be

Rx = Ax + Bx

Rx = 1.80 cm + 4.10 cm

Rx = 5.90 cm

Now Sum of y-component of vectors will be

Ry = Ay + By

Ry = 2.25 cm - 3.50 cm

Ry = -1.25 cm

Sum of both vectors A and B will be

R = A + B

R = (Ax + Bx) i + (Ay + By) j

R = (5.90 i - 1.25 j) cm

Magnitude of R will be

|R| = sqrt (5.90^2 + (-1.25)^2)

|R| = 6.03 cm

Please Upvote.